OpenStudy (anonymous):

Which of the following systems of equations has infinitely many solutions? A. 4x − 7y = 1; 14y = 8x − 3 B. 8x − 3 = 6y; 12y − 6 = 2x C. 4x − 5y = 13; 10y = 8x − 26 D. 5y = -x − 5; -2x + 15y = 10

OpenStudy (anonymous):

you basically solve each set, use addition/subtraction method, until the numbers on the both sides of the = sign are the same. c) 8x-10y=26 -8x+10y=-26 add both together and you get 0=0

OpenStudy (anonymous):

@Katlu95 : Wait thats not the right problem C. im confused